### v1

# Large scale simulations of photosynthetic antenna systems: interplay of cooperativity and disorder

Abstract
Large scale simulations of light-matter interaction in natural photosynthetic antenna complexes containing more than one hundred thousands chlorophyll molecules, comparable with natural size, have been performed. Photosynthetic antenna complexes present in Green sulfur bacteria and Purple bacteria have been analyzed using a radiative non-Hermitian Hamiltonian, well known in the field of quantum optics, instead of the widely used dipole-dipole Frenkel Hamiltonian. This approach allows to study ensembles of emitters beyond the small volume limit (system size much smaller than the absorbed wavelength), where the Frenkel Hamiltonian fails. When analyzed on a large scale, such structures display superradiant states much brighter then their single components. An analysis of the robustness to static disorder and dynamical (thermal) noise, shows that exciton coherence in the whole photosynthetic complex is larger than the coherence found in its parts. This provides evidence that the photosynthetic complex as a whole has a predominant role in sustaining coherences in the system even at room temperature. Our results allow a better understanding of natural photosynthetic antennae and could drive experiments to verify how the response to the electromagnetic radiation depends on the size of the photosynthetic antenna.
keywords:
[…]
1 Introduction
Photosynthesis is a fundamental process able to capture Sun’s energy and convert it into biochemical energy used to drive cellular processes 1. Here we investigate antenna complexes of photosynthetic anaerobic bacteria: the Chlorobium Tepidum Green sulfur bacteria (GSB) and Rhodobacter sphaeroides Purple bacteria (PB) which are among the most efficient photosynthetic complexes in Nature 2. Indeed anaerobic bacteria have the ability to harvest sunlight in deep murky waters, where incident light levels are much reduced beyond the already dilute level on land 3, 4. For instance, Purple bacteria have the ability to exploit extremely weak light sources 3, 4, 5, 6 (less than 10 photons per molecule per second) and some species of Green sulfur bacteria even perform photosynthesis with geothermal radiation from deep-sea hydrothermal vents at about 400∘C 7.
Photosynthetic antenna complexes of anaerobic bacteria 3, 4, 8, 9, 10, 11, 12, 13, 14, 15 are comprised of a network of Bacteriochlorophyll (Bchl) molecules which are typically modelled as two-level systems (2LS) capable of absorbing radiation and transfering the resulting electronic excitation to the reaction center (RC) where charge separation occurs, a process which precedes and drives all other photosynthetic steps. To each 2LS a transition dipole moment (TDM) is associated which determines its coupling with both the electromagnetic field and with other chlorophyll molecules. Owing to the low solar photon density, photosynthetic aggregates operate in the single-excitation regime, meaning that at most one excitation is present in the system at any time. Antenna complexes of photosynthetic bacteria have an internal efficiency of almost (i.e. nearly each photon absorbed produces a charge separation event in the RC) 3, 6.
A possible origin of this incredible ability of bacterial photosynthetic systems to utilize weak sources of incoherent light and funnel the collected energy to specific molecular aggregates could be brought back to the high level of symmetry and hierarchical organization characterizing the antenna complexes of bacterial photosynthetic organisms 16, 17. Photosynthesis in GSB involves chlorophyll pigments tightly packed in light-harvesting systems with cylindrical shapes, known as chlorosomes 1 and shown in Fig. (1). They are able to absorb the sunlight and transfer the electronic excitation to other fundamental units, such as the baseplate and the Fenna-Matthews-Olson (FMO) trimer complex, and finally to the reaction centers (RCs), where the incoming energy is converted into a charged-separated state 3, 6. Purple bacteria offer an alternative organization of the antenna system. The main photosynthetic units of Purple bacteria are the chromatophores ( nm in size) 4 which contain about BChl molecules 18, 19, and are composed of different antenna complexes: LHI, LHII and the RC 4, 20, 21. LHI and LHII complexes show a very well ordered disposition of Bchl molecules which are arranged in ring and S-shaped structures. The energy collected by the chromatophore reaches the RC complex mainly through the LHI complex which surrounds it (see Fig. (2)).
The basic components of the photosynthetic antenna complexes of anaerobic bacteria have been widely studied both theoretically and experimentally in Refs 20, 22, 23, 24, 25, 26, 27. Due to the symmetric arrangement of Bchl molecules these structures display bright (superradiant) and dark (subradiant) states in their single-excitation manifold 28, 3. Bright states are characterized by a giant transition dipole moment (much larger than the single molecule dipole moment), while dark states exhibit a significantly smaller transition dipole moment compared to that of a single molecule. In GSB bright states close to the lower excitonic state have been found in cylindrical Bchl aggregates 29, 26, 30, 31, 32, 33. Additionally, in the PB antenna complex it is well-known that both LHI and LHII complexes display bright states close the lower excitonic state 34, 4, 16, 17, 28.
Nevertheless, the cooperative response to light of the entire photosynthetic antenna complexes has not been theoretically studied thus far. Indeed, whether the arrangement of the Bchl molecules in the entire GSB and PB antenna complexes is capable of supporting collective states brighter than the single sub-units in the complex, is not a trivial question to address. On one side, it is not guaranteed that the symmetry of the entire antenna is capable of supporting a cooperative response larger than its sub-units. Moreover, the most widely used theoretical model, the Frenkel Hamiltonian 35 based on dipole-dipole interaction, becomes ineffective beyond the small volume limit, i.e. when the system size is comparable with the wavelength of the absorbed light. For this reason here we employ a radiative Hamiltonian model, well-established in quantum optics for several decades 12, 29. The radiative Hamiltonian allows us to explore light-matter interaction of photosynthetic antennae beyond the small volume limit, where the Frenkel Hamiltonian cannot be used. The radiative Hamiltonian contains non-Hermitian terms that account for photon losses due to spontaneous emission and accurately describes the effective interaction between molecules. Even if the non-Hermitian part is usually considered in the frame of perturbation theory, this approximation breaks down in large systems when resonances overlap36.
In this study, we thus conduct a large-scale analysis of Green Sulfur Bacteria and Purple Bacteria photosynthetic complexes. Specifically, we examine the GSB chlorosome, which comprises three adjacent concentric cylinders containing more than 105 chlorophyll molecules, see Fig. (1), and the entire chromatophore present in Purple Bacteria, containing approximately 5 000 chlorophyll molecules, see Fig. (2).
We find that cooperative effects in the whole photosynthetic complex are enhanced with respect to its smaller sub-units even when realistic levels of disorder and thermal noise are considered.
2 Models and Methods
The positions and the orientations of the Bchl molecules in both GSB and PB, exhibit a high degree of hierarchical order and symmetry. However, the connection between these structures and their functionality remains an open question. Recent research by some of the authors of this manuscript has demonstrated that the natural arrangement of Bchl molecules in GSB cylindrical structures ensures the presence of cooperative effects even at room temperature 29. In this study, we analyze much larger complexes containing an order of magnitude more Bchl molecules and thus more close to natural sizes.
Below we describe the specific complexes employed in this manuscript and the different Hamiltonian models used to analyze their response to the electromagnetic field. We also describe the theoretical methods used to analyze the robustness of their optical properties to static disorder and thermal noise.
2.1 Antenna complexes in Green Sulfur Bacteria (GSB)
The most common pigments found in the GSB antenna 37, 6 are Bchls c, d, or e, along with carotenoids, and to a lesser extent, Bchls a 38, 24, 39, although Bchls c is the most abundant pigment 40, 41. Pigment organization and orientations of the GSB chlorosome have been studied by using various spectroscopic methods, such as infrared and resonance Raman studies, solid state NMR and cryo-EM, that have revealed that it can be formed by pigments assembling in rod-like (cylindrical) aggregates with lateral lamellae 24, 39, 42, 43, 44, 45, 38. The specific structure are strongly dependent on the growing and the environmental conditions 29, 46, 47. In particular, Chlorobium Tepidum triple mutant antenna complexes present cylindrical structures that can contain between 50 000 to 250 000 Bchlc. These structures typically range in size from to nm in length with widths and depths varying between and nm 37, 48, 49, 50.
Although natural antennae of GSB exhibit considerable variation in size and in the types of Bchl aggregates present (such as lamellae and cylindrical structures), here we model the antenna using only cylindrical structures to capture the primary natural features. The cylindrical structures in GSB are made of concentric single wall cylinders, which are observed to lie adjacently to each other on the baseplate.
The model that we consider is shown in Fig. (1). It is inspired by the antenna complexes found in the Green sulfur bacterium C. Tepidum bchQRU triple mutant type, whose chlorosomes exhibit a much more regular geometry with respect to the wild type. Nevertheless, both types exhibit very similar optical properties and demonstrate similar cooperative effects 29, 51. Below, we describe, in order of hierarchical complexity, the geometry of three structures utilized in this study to model the cooperative effects occurring in the antenna systems of photosynthetic GSB.
Complex A - single cylinder. In Fig.(1 A) we present the model utilized for a single cylinder in the bchQRU triple mutant chlorosome. The cylinder comprises a stack of tightly-packed rings of Bchl c molecules. Each molecule is treated as a dipole with a squared dipole moment , possessing a well defined orientation and position in space 29. Each ring has a radius and contains Bchl molecules separated by a nearest neighbour distance . The rings are equally separated by a vertical distance 29, 51, 38. In our simulations the single cylinder complex can contain up to 10 800 Bchl c molecules, resulting in a maximum length , corresponding to a stack of rings. Further details about the geometry of the bchQRU triple mutant concerning the positions and orientations can be found in Ref. 29.
Complex B - four concentric cylinders. Chlorosomes of GSB display a complex arrangement of molecules on a multi-wall structure 29, 38, 6, 39, 52. Panel B of Fig. (1) depicts the model we developed, wherein Bchl c molecules form four concentric cylinders. On each cylinder, the dipoles positions and orientations are consistent with those found in the single cylinder model of the bchQRU triple mutant, as explained in 29. The innermost cylinder has a radius nm, while the distance from wall to wall is nm 47, 29.
Complex C - three adjacent concentric cylinders. In panel C of Fig. (1) we show a more elaborate model formed by three adjacent structures, each comprising four concentric cylinders. The wall-to-wall distance between adjacent cylinders is set to nm, a realistic value according to Ref. 24. The maximum length considered here is nm, corresponding to vertical layers in each cylindrical structure. The largest system considered in our simulations contains 132 840 Bchl c molecules, comparable with natural sizes. The positions and orientations of dipoles on each tubular wall are the same as in Complex A and B.
2.2 Antenna complexes in Purple bacteria (PB).
In Purple bacteria Rhodobacter Sphaeroides, the light-harvesting complexes are organized in thousands of spherical membrane-embedded protrusions, called chromatophores which typically measure around nm in diameter. Depending on the light conditions, Purple bacteria can contain from up to 2 500 chromatophores made of Bchl molecules each 4. The major pigments in the chromatophores are the Bchl a 18, 19 which are organized on different complexes inside the chromatophores. Here we consider a portion of the chromatophore (complex A) and the whole chromatophore (complex B):
Complex A - light-harvesting system subunit. The RC is directly surrounded by the so-called light harvesting complex I (LHI). LHI is in turn surrounded by several smaller light harvesting complexes, the LHII complexes 2, 53, 4. As shown in panel A of Fig. (2) two RCs are surrounded by the light-harvesting complex LHI B875, which is an S-shaped structure with 56 BChl molecules. The LHI aggregate is a J-aggregate with two superradiant states at 875 nm that are polarized in the ring plane and close to the lowest excitonic state. The LHI aggregate is surrounded by ten LHII rings, each featuring the B850 ring, a J-aggregate composed of 18 BChl molecules with two superradiant states at 850 nm, and the B800 ring composed of 9 BChl molecules with the main absorption peak at 800 nm. This hierarchical structure is able to absorb photons at different frequencies and to funnel the collected energy to the RC 54, 34, 4, 17, 16. In our simulation we model the RC as an aggregate of four BChls: two forming the tightly-coupled special pair and two accessory ones 17. Even if other molecules are present in the RC, these four Bchl molecules are the most relevant for the interaction with the electromagnetic field 55.
Complex B: the chromatophore. We model the whole 3D structure of the chromatophore using data from Ref. 34 collected by AFM, cryo-EM, X-ray crystallography and NMR measurements. In panel B of Fig. (2) our model of the chromatophore is shown with different substructures: i) LHII (blue rings); ii) LHI (orange structures); iii) RC, placed at the center of the LHI structures. Chromatophores, see Fig. (2 B), are connected to the cell membrane at their south pole, therefore the southern polar region is left empty (no molecular aggregates are placed there) to allow the contact with the cell membrane 56. The chromatophore spherical vesicles are reproduced by mapping small planar regions onto spherical ones containing the LHI-RCs or an LHII complex with the area-preserving inverse-Mollweide transformations, see Fig. (LABEL:PBprojection) in LABEL:AppA of the Supporting Information, used also in Ref. 34, 57, 58. While in nature the ratio between LHII and LHI-RC depends on illumination and other conditions, here we consider a vesicle model that contains S-shaped LHI-RCs and LHII complexes. The overall Bchl/RC ratio used in our simulation () is within the natural range for these vesicles, which is between and and corresponds to low-light growth conditions 34 where these systems show the largest efficiency. The data for the LHI, LHII and RC are taken from Ref. 17, 4. Even if it would be more realistic to add random rotations of LHII rings around the axis joining the center of the vesicle and the center of the LHII rings, we checked that the effect of such random rotations is small, see Fig. (LABEL:PBrotationLHII) of LABEL:AppE, and thus we did not include such rotations in our simulations.
2.3 The radiative Hamiltonian and the dipole approximation
Since photosynthetic antennae operate under natural sunlight, which is very dilute, the single-excitation approximation can be used, so that only states containing a single excitation have been considered. Choosing the basis states in the single-excitation manifold, where represents a state in which the molecule is excited while all the others are in the ground state, the systems can be described through a radiative non-Hermitian Hamiltonian which takes into account the interaction between the molecules mediated by the electromagnetic field (EMF) 13, 12, 59, 12, 60. The radiative Hamiltonian reads:
where is the excitation energy of single emitter (Bchl molecule in our case). The terms and have a diagonal part given by:
with being the transition dipole moment (TDM) and , where is the wavelength associated with the molecular transition. The off-diagonal part () is given by
where is the unit dipole moment of the site and is the unit vector joining the and the sites. See LABEL:AppB of the Supporting Information for the parameters we used for GSB and PB in the Hamiltonian.
Diagonalizing the Hamiltonian (1) we obtain the complex eigenvalues
where is the radiative decay of the eigenstate. In general differs from the radiative decay of the single molecule . In particular, when the ratio we will talk about a “superradiant state” (SRS) or bright state, otherwise when the state is called “subradiant” or dark. In other words, a SRS can radiate much faster than a single molecule, while a subradiant one radiates at a rate much slower than the single molecule radiative decay.
If the non-Hermitian term can be considered a small perturbation, the optical absorption of an eigenstate can be estimated in terms of its dipole strength, computed using only the Hermitian part of the Hamiltonian (1). Denoting the eigenstate of the Hermitian part of the Hamiltonian (1) with , we can expand it on the site basis, so that
To each basis state , a dipole moment is associated, corresponding to the TDM of the molecule. If is the total number of molecules, then we will express the TDM associated with the eigenstate as follows:
The dipole strength of the eigenstate is defined by (note that due to normalization ).
The non-Hermitian part of the radiative Hamiltonian (1) can be treated as a perturbation whenever the decay widths are much smaller than the mean level spacing computed using the real part of the complex eigenvalues. When this criterion, known as resonance overlap criterion 36, is valid, one can exclusively utilize the Hermitian part of the Hamiltonian. This reduction in complexity accelerates calculations. The Hermitian part of the Hamiltonian (1) is defined as follows:
where is given in Eq. (3). Note that when the resonances do not overlap we have .
Finally, we note that when resonances do not overlap and the system size is much smaller than (i.e. when ), the Hermitian part of the radiative Hamiltonian reduces to the standard dipole-dipole Frenkel Hamiltonian:
In the following, we will analyzed all complexes using the three different Hamiltonian models introduced in this section, namely:
1.
NHH: non-Hermitian radiative Hamiltonian Eq. (1).
2.
HH: Hermitian Hamiltonian Eq. (7) valid under the non-overlapping resonance criterion.
3.
DH: Dipole Hamiltonian Eq. (8) valid under the non-overlapping resonance criterion and when the system size is small compared to the wavelength associated with the optical transition of the molecules.
2.4 Robustness to thermal noise and static disorder
Given a quantum state specified by the density matrix it is possible to define its coherence length in the single-excitation manifold defined by the basis states 61, 62, 29 in the following way:
in Eq. (9) measures how much a single excitation is spread coherently over the molecules composing the aggregate. To give an idea of its physical meaning let us consider three different simple cases:
a pure localized state, ; then it is easy to see that the coherence length defined in Eq. (9) is given by . This case represents the minimal value that can get.
A completely delocalized mixed state characterized by the density matrix
(10) In this case we have . This state is maximally delocalized in the basis, but it is completely incoherent.
Lastly we consider the fully delocalized coherent state: . In this case we have . Note that any pure state with constant amplitude over the sites and arbitrary phases would give the same result.
It is easy to see that . The closer is to , the higher a coherent delocalization can be assigned to our state. In the same way indicates an incoherent localized state. States characterized by have a little ambiguity (since both localization and coherence are measured on the same length scale).
For all models we have computed the thermal coherence length at room temperature ( K), defined for a state at the canonical equilibrium and whose matrix elements are given by:
where .
A very important question to be answered is how much the symmetrical arrangement of the molecules that produces superradiance,
is also able to produce a large thermal coherence length at room temperature. Note that even if we consider the coherence length at thermal equilibrium, this does not mean that out-of-equilibrium processes are not important in molecular nanotubes. Indeed in Ref. 63 strong evidence of ultra-fast transport in natural structures with transfer times less than 100 fs have been discussed. Nevertheless thermal equilibrium can be considered as a worst case scenario for coherences. For this reason assuming thermal equilibrium can be considered a good starting point to assess the structural robustness of quantum coherence to thermal noise.
In the following we calculate the coherence length according to Eq. (9), using a thermal density matrix as in Eq. (11).
Natural complexes are not only affected by thermal noise but also by other sources of disorder due to the fluctuations in the local environment. In order to analyze the robustness to this kind of disorder, we have considered time-independent and space-dependent fluctuations of the excitation energy of the molecules in the aggregates. Specifically, we consider energy fluctuations which are uniformly distributed around the excitation energy of the molecules , between and , where represents the disorder strength. It is known that static disorder induces eigenmodes localization 64 and quenching of superradiance 65.
3 Numerical results
In this section we show the numerical results obtained for the antenna complexes of GSB and PB.
3.1 Green Sulfur Bacteria complexes
Here, the complexes A, B, C of GSB described in Sec. 2.1 have been analyzed. In Fig. (3) the spectrum of the three complexes is shown for the maximal length considered, nm, corresponding to cylinders with 180 layers. Thus, for the three complexes we have (complex A), (complex B), (complex C) chlorophyll molecules. Panels (A-C) in Fig. (3) show that the structure of all complexes allow for the emergence of red-shifted superradiant states, close to the lowest excitonic states. Moreover the amount of maximal superradiance () increases with the system size, see Fig. (4). The largest amount of superradiance is present in complex C which is the largest one and the closest to the natural size and structure of GSB antennae. Note that the fact that in larger aggregates superradiance is enhanced is far from being trivial. Indeed, if we would increase the system size by adding the molecules in the same positions but with randomized dipole directions, no superradiant enhancement would be present, as shown in LABEL:AppF-GSB of the Supporting Information.
By comparing the three Hamiltonians, see also discussion in LABEL:AppC, we note that for complex A, all of them give a good description of the superradiant states. For complex B a deviation between DH and the other two models (NHH and HH) is observed, while in complex C the three Hamiltonian models give different results. This shows that using the most accurate Hamiltonian, which is the NHH model, is essential to describe large photosynthetic antennae. The same behavior is also shown in Fig. (4). In each panel of Fig. (4) we show the maximal decay width for each complex as a function the complex length. In each panel the maximal decay width has been computed using three different Hamiltonian models: DH, HH, NHH. For the DH and HH model we computed the maximal decay width using the dipole strength, while for the NHH model the maximal decay width has been computed using the imaginary part of the complex eigenvalues of the radiative non-Hermitian Hamiltonian. As one can see in the largest complex Fig. (4 panel C) the three Hamiltonians give very different results.
The origin of the difference between the NHH and the other Hamiltonian models can be explained by the overlapping resonance criterion 36, see panel D in Fig. (3). Indeed, when the decay width becomes comparable with the eigenmodes mean level spacing, the non-Hermitian part of the Hamiltonian cannot be treated perturbatively and the dipole strength Eq. (6) does not describe anymore the decay widths of the eigenmodes.
In order to analyze the robustness of natural models to static disorder and thermal noise we have chosen two figures of merit: the maximal dipole strength Eq. (6) and the thermal coherence length Eq. (9). Both quantities have been studied as a function of the static disorder strength .
Fig. (5 A,B) shows the maximal dipole strength and the thermal coherence length for the case of a single cylinder (complex A of the GSB) for different cylinder lengths. For complex A we used the HH model to compute both figures of merits since the results obtained with the latter model do not differ substantially from those obtained with the NHH model while giving a substantial computational advantage.
In Fig. (5 A,B), one can see that both quantities increase with the system size, even if the thermal coherence length tends to saturate. Moreover, both figures of merits show that single cylinders are very robust to disorder and thermal noise. Indeed, their maximal value (for ) remains mostly unchanged up to values of the disorder strength comparable with the thermal energy at room temperature ( for K). Note that, in natural systems, the static disorder strength is usually of the same order of magnitude of the thermal energy. In Fig. (5 C,D) the maximal dipole strength (panel C) and the thermal coherence length (panel D) are shown at and for different cylinder length containing different number of chlorophyll molecules. Interestingly the value of both figure of merits is the same for the two values of disorder strength considered, showing their extreme robustness to disorder. Moreover the maximal dipole strength (panel C) increases linearly with which shows that such structures are extremely effective at preserving cooperative effects as a function of the system size. Indeed single excitation superradiance cannot increase faster than . The thermal coherence also increases with the system size (panel D), even if it shows a tendency to saturate as increases.
We now consider in Fig. (6), the three different complexes for a fixed length corresponding to 100 layers ( nm), smaller than the maximal system length considered in Fig. (3,4). Also in this case we used the HH model to compute both figures of merit since for this length scale the HH model is quite accurate for all complexes, see discussion in LABEL:AppC of the Supporting Information. While the maximal dipole strength decreases with disorder Fig. (6 A), it is still much larger than one even for . On the other hand, the thermal coherence length shows a large robustness to disorder, Fig. (6 B) up to . In Fig. (6 C,D) the maximal dipole strength (panel C) and the thermal coherence length (panel D) are shown at and as a function of the number of chlorophyll molecules present in the three different complexes A,B,C. The most important feature of both figures of merit is that they grow as a larger portion of the photosynthetic antenna is considered, showing that the structure of the GSB photosynthetic antenna as a whole is able to support excitonic coherences. Note that in presence of disorder, the maximal dipole strength grows slower with compared with the case of zero disorder (panel C). On the hand, the value of the coherence length is the same for the two values of disorder strength considered, showing its robustness to disorder (panel D), but as increases it shows a tendency to saturate.
3.2 Purple bacteria Complexes
Here we analyze the whole chromatophore of the Purple bacteria antenna complex (complex B), see Fig. (2 B) containing 4 113 chlorophyll molecules. As a comparison we will also analyze complex A, see Fig. (2 A).
Fig. (7) shows the spectra of the two complexes obtained by diagonalizing the three Hamiltonians models (HH, NHH and DH). Panels (A-B) of Fig. (7) demonstrate that both PB complexes allow the emergence of red-shifted superradiant states. Furthermore, the amount of superradiance is larger in the larger complex, the chromatophore. It is interesting to note that all Hamiltonian models (HH, NHH and DH) give very similar results in both the complexes considered. The HH model and the NHH model give similar results since, as it is shown in panel C of Fig. (7), the decay widths are always smaller than the eigenmodes mean level spacing, so that the non-Hermitian part of the Hamiltonian can be treated perturbatively. The DH model is also a good approximation even for the whole chromatophore since so that we are in the small volume limit.
Finally, we analyze the robustness to static disorder and thermal noise using the average maximal dipole strength (6) and the average thermal coherence length (9) as figures of merit. Both of them have been computed using the HH model.
In Fig. (8) the average maximal dipole strength and thermal coherence length for both complexes A and B are shown as a function of the disorder strength rescaled over the thermal energy at room temperature, K. For sake of comparison, we also add the data for the S-shaped LHI (green symbols), indicated by the cyan structure in Fig. (2 A), which is smaller than complex A and complex B.
As for the GSB complexes, the maximal dipole strength decreases with disorder Fig. (8 A), even if it is still much larger than one for . On the other hand, the thermal coherence length shows a large robustness to disorder, Fig. (8 B) up to . In Fig. (8 C,D) the maximal dipole strength (panel C) and the thermal coherence length (panel D) are shown at and as a function of the number of chlorophyll molecules contained in the S-shaped LHI and complexes A and B.
Both figures of merit grow as a larger portion of the photosynthetic antenna is considered. Note that in presence of disorder, the maximal dipole strength grows slower with compared with the case of zero disorder (panel C). On the hand, the value of the coherence length is the same for the two values of disorder strength considered, showing its robustness to disorder (panel D). Nevertheless, as increases, it shows a tendency to saturate.
4 Conclusions
In this manuscript large scale simulations of the whole antenna complexes of Green sulfur bacteria and Purple bacteria have been performed using a radiative non-Hermitian Hamiltonian, which, at variance with the standard Frenkel dipole Hamiltonian, is valid beyond the small volume limit. The largest Green sulfur antenna complex considered was composed of 132 840 Bchlc molecules, arranged in three adjacent cylinders, each composed of 4 concentric cylinders of different radii. The size of the largest complex was , comparable with the size of natural antenna complexes. For the Purple bacteria the whole chromatophore was considered, which is a spherical structure of of diameter and composed of 4 113 Bchla molecules.
While smaller portions of these complexes have been widely analyzed in literature and superradiance has been found both theoretically and experimentally, an open question was whether superradiant effects will be enhanced in the whole complex. Our analysis has shown that the maximal superradiant decay width in the largest Green sulfur bacteria complex is ten times larger than in a single wall cylinder of the same length (see Fig. (6) panel A), while in the Purple bacteria chromatophore, it is about six times larger than in the single LHI complex and about ten times larger than the single LHII complex (see Fig. (8) panels A and B). This proves that the bare structure of the whole antenna complexes is able to support an enhanced superradiant response. This is not trivial since superradiance critically depends on the arrangement of the emitters and it can be easily quenched. Also the effect of realistic values of static disorder and thermal noise have been investigated. The cooperative effects in large complexes have been found to be robust even with disorder and noise levels comparable with ambient conditions. In particular, the thermal coherence length in the largest Green sulfur bacteria complex is between two and three times larger than in a single wall cylinder of the same length, see Fig. (6 panel B) while in the Purple bacteria chromatophore the thermal coherence length is approximately two times larger than in the single (S-shaped) LHI complex.
In analyzing different antenna complexes, we compared three distinct Hamiltonian models: the Dipole Hamiltonian, applicable in the small volume limit; the Hermitian Hamiltonian, valid when the decay widths are significantly smaller than the mean level spacing; and the full radiative non-Hermitian Hamiltonian model. For Purple bacteria chromatophores, all models yield consistent results due to the system’s small size (e.g., 60 nm) relative to the absorbed wavelength () and the limited number of Bchl molecules present. However, for the largest complexes considered in Green sulfur bacteria, only the non-Hermitian Hamiltonian model proved effective in providing reliable results.
The presence of very bright (superradiant) states in the whole photosynthetic antenna could shed new light in understanding the functionality of such complexes. In particular, the presence of superradiance and subradiance plays an important role for the efficiency of the energy transfer in these aggregates 66, 17. Let us also remark that the system dynamics strongly depends on the choice of the initial state which is mainly determined by the most superradiant states, see LABEL:AppD for a more detailed discussion about the delocalization of superradiant states in PB and GSB antennae.
Our theoretical findings could inspire experimental validation of the extent of cooperative response in natural complexes predicted in this manuscript. To address these experiments, spectroscopy methods can be employed to characterize the size dependence of the optical response, comparing different portions of the photosynthetic antennae extracted from Green and Purple bacteria, with the whole photosynthetic structures. In particular, the radiative decay widths obtained through time resolved fluorescence spectroscopy and the quantum yield obtained from spectrofluorimetric measurements conducted at different temperatures (from to room temperature) could provide a clear signature of cooperative effects present in large photosynthetic aggregates.
The cooperative properties, akin to those observed in photosynthetic aggregates, are inspiring numerous proposals for engineering artificial devices for light-harvesting and clean energy production 67, 68, 69, 70, 71, 72, 73, 74. Thus, our study could significantly impact the development of quantum devices for photon detection and light harvesting.
We acknowledge useful discussion with M. Trotta and Anna Painelli. We also acknowlege P. Bolpagni and F. Airoldi for their valuable contribution in the initial stage of this work. We acknowlege L. Pattelli for managing the computer code on one of the LENS clusters and G. Mazzamuto for his help to optimize numerical simulations. E. Pini is kindly acknowledged for help with the simulations. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute of theoretical Physics (KITP). G. L. C. acknowledges financial support from PNRR MUR project PE0000023-NQSTI. The authors would like to acknowledge the High Performance Computing Center of the University of Strasbourg for supporting this work by providing scientific support and access to computing resources. Part of the computing resources were funded by the Equipex Equip@Meso project (Programme Investissements d’Avenir) and the CPER Alsacalcul/Big Data. The authors also thank the Institute of Physics BUAP Mexico for allowing the access to their computational resources.